OpenStudy (anonymous):

hola so I'm doing a few path problems and I'm getting behind in math I have 4 questions I really need help with. anybody up for the task? Thanks!

OpenStudy (anonymous):

*math any help would be awesome

OpenStudy (anonymous):

what do you need help with?

OpenStudy (anonymous):

It's a little complicated how do you upload pics so you can see my questions clearly?

OpenStudy (anonymous):

where you type the reply theres little buttons under where youre typing press attach file and choose your pictures.

OpenStudy (anonymous):

\[-9\sqrt{28a ^{2}} \times \frac{ 1 }{ 3 } \sqrt{63a}\]

OpenStudy (anonymous):

hello??